Miguel is 46 years old. He is 4 years older than thrice his son’s age. Find the age of his son.
Answer by sravancv(7)

Given
Miguel is 46 years old.
Let his son age be "x"
also given than he is 4 years older than thrice of his son's age
so
3x+4 = 46
3x=46-4
3x=42
x=42/3
x=14
so his son age is 14 years.